OpenStudy (anonymous):

is 5/8 greater than, less than or equal to 16/21? **Please help ;3**

OpenStudy (anonymous):

5/8 * 21/21 = 105/21*8 16/21 * 8/8 = 128/21*8

OpenStudy (anonymous):

thank you! c:

OpenStudy (anonymous):

np :D